OpenStudy (anonymous):

im_{x rightarrow 1} sqrt[n]{x}-sqrt[m]{x}div sqrt[p]{x}-1

OpenStudy (anonymous):

solve \[\lim_{x \rightarrow 1}\sqrt[n]{x}-\sqrt[m]{x}\div \sqrt[p]{x}-1\] ?